% Generated by IEEEtran.bst, version: 1.14 (2015/08/26)
\begin{thebibliography}{10}
\providecommand{\url}[1]{#1}
\csname url@samestyle\endcsname
\providecommand{\newblock}{\relax}
\providecommand{\bibinfo}[2]{#2}
\providecommand{\BIBentrySTDinterwordspacing}{\spaceskip=0pt\relax}
\providecommand{\BIBentryALTinterwordstretchfactor}{4}
\providecommand{\BIBentryALTinterwordspacing}{\spaceskip=\fontdimen2\font plus
\BIBentryALTinterwordstretchfactor\fontdimen3\font minus
  \fontdimen4\font\relax}
\providecommand{\BIBforeignlanguage}[2]{{%
\expandafter\ifx\csname l@#1\endcsname\relax
\typeout{** WARNING: IEEEtran.bst: No hyphenation pattern has been}%
\typeout{** loaded for the language `#1'. Using the pattern for}%
\typeout{** the default language instead.}%
\else
\language=\csname l@#1\endcsname
\fi
#2}}
\providecommand{\BIBdecl}{\relax}
\BIBdecl

\bibitem{heldens2020landscape}
S.~Heldens, P.~Hijma, B.~V. Werkhoven, J.~Maassen, A.~S. Belloum, and R.~V.
  Van~Nieuwpoort, ``The landscape of exascale research: A data-driven
  literature analysis,'' \emph{ACM Computing Surveys (CSUR)}, vol.~53, no.~2,
  pp. 1--43, 2020.

\bibitem{hoekstra2019multiscale}
A.~G. Hoekstra, B.~Chopard, D.~Coster, S.~Portegies~Zwart, and P.~V. Coveney,
  ``Multiscale computing for science and engineering in the era of exascale
  performance,'' \emph{Philosophical Transactions of the Royal Society A}, vol.
  377, no. 2142, p. 20180144, 2019.

\bibitem{bernholdt2020survey}
D.~E. Bernholdt, S.~Boehm, G.~Bosilca, M.~Gorentla~Venkata, R.~E. Grant,
  T.~Naughton, H.~P. Pritchard, M.~Schulz, and G.~R. Vallee, ``A survey of mpi
  usage in the us exascale computing project,'' \emph{Concurrency and
  Computation: Practice and Experience}, vol.~32, no.~3, p. e4851, 2020.

\bibitem{schneider2022exascale}
D.~Schneider, ``The exascale era is upon us: The frontier supercomputer may be
  the first to reach 1,000,000,000,000,000,000 operations per second,''
  \emph{IEEE Spectrum}, vol.~59, no.~1, pp. 34--35, 2022.

\bibitem{alexander2020exascale}
F.~Alexander, A.~Almgren, J.~Bell, A.~Bhattacharjee, J.~Chen, P.~Colella,
  D.~Daniel, J.~DeSlippe, L.~Diachin, E.~Draeger \emph{et~al.}, ``Exascale
  applications: skin in the game,'' \emph{Philosophical Transactions of the
  Royal Society A}, vol. 378, no. 2166, p. 20190056, 2020.

\bibitem{murphy2007effects}
R.~Murphy, ``On the effects of memory latency and bandwidth on supercomputer
  application performance,'' in \emph{2007 IEEE 10th International Symposium on
  Workload Characterization}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE,
  2007, pp. 35--43.

\bibitem{pfister2001introduction}
G.~F. Pfister, ``An introduction to the infiniband architecture,'' \emph{High
  performance mass storage and parallel I/O}, vol.~42, no. 617-632, p. 102,
  2001.

\bibitem{grun2010introduction}
P.~Grun, ``Introduction to infiniband for end users,'' \emph{White paper,
  InfiniBand Trade Association}, vol.~55, 2010.

\bibitem{ajima2022optical}
Y.~Ajima, ``Optical connection of top-level supercomputers: current status and
  future expectations,'' \emph{Optical Interconnects XXII}, vol. 12007, pp.
  11--19, 2022.

\bibitem{lu2022survey}
P.-J. Lu, M.-C. Lai, and J.-S. Chang, ``A survey of high-performance
  interconnection networks in high-performance computer systems,''
  \emph{Electronics}, vol.~11, no.~9, p. 1369, 2022.

\bibitem{vergara2019scaling}
V.~G. Vergara~Larrea, W.~Joubert, M.~J. Brim, R.~D. Budiardja, D.~Maxwell,
  M.~Ezell, C.~Zimmer, S.~Boehm, W.~Elwasif, S.~Oral \emph{et~al.}, ``Scaling
  the summit: deploying the world’s fastest supercomputer,'' in
  \emph{International Conference on High Performance Computing}.\hskip 1em plus
  0.5em minus 0.4em\relax Springer, 2019, pp. 330--351.

\bibitem{parashar2013cloud}
M.~Parashar, M.~AbdelBaky, I.~Rodero, and A.~Devarakonda, ``Cloud paradigms and
  practices for computational and data-enabled science and engineering,''
  \emph{Computing in Science \& Engineering}, vol.~15, no.~4, pp. 10--18, 2013.

\bibitem{aws2}
AWS, ``Amazon ec2 hpc6a instances,''
  \url{https://aws.amazon.com/ec2/instance-types/hpc6/}, January 10, 2022,
  [Online].

\bibitem{aws3}
AWS, ``Amazon ec2 hpc6a instance optimized for high performance computing,''
  \url{https://aws.amazon.com/blogs/aws/new-amazon-ec2-hpc6a-instance-optimized-for-high-performance-computing/},
  January 10, 2022, [Online].

\bibitem{gcp4}
G.~C.~A. Center, ``Compute-optimized machine family,''
  \url{https://cloud.google.com/compute/docs/compute-optimized-machines },
  November 30, 2021, [Online].

\bibitem{walkup2022best}
R.~Walkup, S.~R. Seelam, and S.~Wen, ``Best practices for hpc workloads on
  public cloud platforms: A guide for computational scientists to use public
  cloud for hpc workloads,'' in \emph{Proceedings of the 2022 ACM/SPEC on
  International Conference on Performance Engineering}, 2022, pp. 29--35.

\bibitem{fernandez2021evaluation}
A.~Fernandez, ``Evaluation of the performance of tightly coupled parallel
  solvers and mpi communications in iaas from the public cloud,'' \emph{IEEE
  Transactions on Cloud Computing}, 2021.

\bibitem{aws1}
J.~Barr, ``Elastic fabric adapter (efa) for tightly-coupled hpc workloads,''
  \url{https://aws.amazon.com/blogs/aws/now-available-elastic-fabric-adapter-efa-for-tightly-coupled-hpc-workloads/},
  April 29, 2019, [Online].

\bibitem{gcp1}
D.~Phanekham, ``What a trip! measuring network latency in the cloud,''
  \url{https://cloud.google.com/blog/products/networking/using-netperf-and-ping-to-measure-network-latency
  }, June 16, 2020, [Online].

\bibitem{gcp2}
J.~Glenn, ``Network latency and throughput— measuring those metrics for cloud
  resources,''
  \url{https://faun.pub/network-latency-and-throughput-measuring-those-metrics-for-cloud-resources-cf8b919b6a98},
  July 28, 2020, [Online].

\bibitem{gcp3}
G.~C.~A. Center, ``Best practices for running tightly coupled hpc applications
  on compute engine,''
  \url{https://cloud.google.com/architecture/best-practices-for-using-mpi-on-compute-engine
  }, November 30, 2021, [Online].

\bibitem{phanekham2020measuring}
D.~Phanekham, M.~Zaber, S.~Nair, S.~Deitz, R.~Jones, M.~Chalasani, and
  M.~Truty, ``Measuring cloud network performance with perfkit benchmarker,''
  \emph{Technical White Paper. Google Cloud}, 2020.

\bibitem{aws6}
J.~Beswick, ``Using shared memory for low-latency, intra-node communication in
  aws batch,''
  \url{https://aws.amazon.com/blogs/compute/using-shared-memory-for-low-latency-intra-node-communication-in-aws-batch/},
  May 6, 2021, [Online].

\bibitem{zivanovic2016large}
D.~Zivanovic, M.~Radulovic, G.~Llort, D.~Zaragoza, J.~Strassburg, P.~M.
  Carpenter, P.~Radojkovi{\'c}, and E.~Ayguad{\'e}, ``Large-memory nodes for
  energy efficient high-performance computing,'' in \emph{Proceedings of the
  Second International Symposium on Memory Systems}, 2016, pp. 3--9.

\bibitem{kashyap2022nvme}
A.~Kashyap and X.~Lu, ``Nvme-oaf: Towards adaptive nvme-of for io-intensive
  workloads on hpc cloud,'' in \emph{Proceedings of the 31st International
  Symposium on High-Performance Parallel and Distributed Computing}, 2022, pp.
  56--70.

\bibitem{guz2018performance}
Z.~Guz, H.~Li, A.~Shayesteh, and V.~Balakrishnan, ``Performance
  characterization of nvme-over-fabrics storage disaggregation,'' \emph{ACM
  Transactions on Storage (TOS)}, vol.~14, no.~4, pp. 1--18, 2018.

\bibitem{gupta2011evaluation}
A.~Gupta and D.~Milojicic, ``Evaluation of hpc applications on cloud,'' in
  \emph{2011 Sixth Open Cirrus Summit}.\hskip 1em plus 0.5em minus 0.4em\relax
  IEEE, 2011, pp. 22--26.

\bibitem{gcp5}
G.~C.~A. Center, ``About local ssds,''
  \url{https://cloud.google.com/compute/docs/disks/local-ssd }, July 28, 2022,
  [Online].

\bibitem{aws4}
V.~Boya, ``Amazon ec2 r6id instances with nvme local instance storage of up to
  7.6 tb,''
  \url{https://aws.amazon.com/blogs/aws/new-amazon-ec2-r6id-instances/ }, June
  9, 2022, [Online].

\bibitem{netto2018hpc}
M.~A. Netto, R.~N. Calheiros, E.~R. Rodrigues, R.~L. Cunha, and R.~Buyya, ``Hpc
  cloud for scientific and business applications: taxonomy, vision, and
  research challenges,'' \emph{ACM Computing Surveys (CSUR)}, vol.~51, no.~1,
  pp. 1--29, 2018.

\bibitem{top500}
top500, ``Top500, november 2021,''
  \url{https://www.top500.org/lists/top500/2021/11/ }, November, 2021,
  [Online].

\bibitem{napper2009can}
J.~Napper and P.~Bientinesi, ``Can cloud computing reach the top500?'' in
  \emph{Proceedings of the combined workshops on UnConventional high
  performance computing workshop plus memory access workshop}, 2009, pp.
  17--20.

\bibitem{denning2016exponential}
P.~J. Denning and T.~G. Lewis, ``Exponential laws of computing growth,''
  \emph{Communications of the ACM}, vol.~60, no.~1, pp. 54--65, 2016.

\bibitem{jing2020cache}
C.~Jing and H.~Zhu, ``Cache and memory,'' in \emph{Data Plane Development Kit
  (DPDK)}.\hskip 1em plus 0.5em minus 0.4em\relax CRC Press, 2020, pp. 31--49.

\bibitem{aurora}
``Aurora | argonne leadership computing facility,''
  \url{https://alcf.anl.gov/aurora}, 2021, [Online].

\bibitem{luttgau2018survey}
J.~L{\"u}ttgau, M.~Kuhn, K.~Duwe, Y.~Alforov, E.~Betke, J.~Kunkel, and
  T.~Ludwig, ``Survey of storage systems for high-performance computing,''
  \emph{Supercomputing Frontiers and Innovations}, vol.~5, no.~1, pp. 31--58,
  2018.

\bibitem{harrod2012journey}
W.~Harrod, ``A journey to exascale computing,'' in \emph{2012 SC Companion:
  High Performance Computing, Networking Storage and Analysis}.\hskip 1em plus
  0.5em minus 0.4em\relax IEEE, 2012, pp. 1702--1730.

\bibitem{fan2001memory}
X.~Fan, C.~S. Ellis, and A.~R. Lebeck, ``Memory controller policies for dram
  power management,'' in \emph{ISLPED'01: Proceedings of the 2001 International
  Symposium on Low Power Electronics and Design (IEEE Cat. No.
  01TH8581)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2001, pp. 129--134.

\bibitem{oren2016memory}
G.~Oren, L.~Barenboim, and L.~Amar, ``Memory-aware management for heterogeneous
  main memory using an optimization of the aging paging algorithm,'' in
  \emph{2016 45th International Conference on Parallel Processing Workshops
  (ICPPW)}.\hskip 1em plus 0.5em minus 0.4em\relax IEEE, 2016, pp. 98--105.

\bibitem{bergman2008exascale}
K.~Bergman, S.~Borkar, D.~Campbell, W.~Carlson, W.~Dally, M.~Denneau,
  P.~Franzon, W.~Harrod, K.~Hill, J.~Hiller \emph{et~al.}, ``Exascale computing
  study: Technology challenges in achieving exascale systems,'' \emph{Defense
  Advanced Research Projects Agency Information Processing Techniques Office
  (DARPA IPTO), Tech. Rep}, vol.~15, 2008.

\bibitem{Argonne1}
A.~N. Laboratory, ``Aurora,'' \url{https://www.alcf.anl.gov/aurora/}, 2022.

\bibitem{intel1}
Intel, ``Solve data challenges with memory tiering,''
  \url{https://www.intel.com/content/www/us/en/products/docs/memory-storage/optane-persistent-memory/memory-tiering-improving-data-management-paper.html
  }, April, 2022, [Online].

\bibitem{hirofuchi2020prompt}
T.~Hirofuchi and R.~Takano, ``A prompt report on the performance of intel
  optane dc persistent memory module,'' \emph{IEICE TRANSACTIONS on Information
  and Systems}, vol. 103, no.~5, pp. 1168--1172, 2020.

\bibitem{hirofuchi2019preliminary}
T.~Hirofuchi and R.~Takano, ``The preliminary evaluation of a hypervisor-based
  virtualization mechanism for intel optane dc persistent memory module,''
  \emph{arXiv preprint arXiv:1907.12014}, 2019.

\bibitem{patil2021nvm}
O.~Patil, L.~Ionkov, J.~Lee, F.~Mueller, and M.~Lang, ``Nvm-based energy and
  cost efficient hpc clusters,'' 2021.

\bibitem{fridman2021assessing}
Y.~Fridman, Y.~Snir, M.~Rusanovsky, K.~Zvi, H.~Levin, D.~Hendler, H.~Attiya,
  and G.~Oren, ``Assessing the use cases of persistent memory in
  high-performance scientific computing,'' in \emph{2021 IEEE/ACM 11th Workshop
  on Fault Tolerance for HPC at eXtreme Scale (FTXS)}.\hskip 1em plus 0.5em
  minus 0.4em\relax IEEE, 2021, pp. 11--20.

\bibitem{hennecke2020daos}
M.~Hennecke, ``Daos: A scale-out high performance storage stack for storage
  class memory,'' \emph{Supercomputing frontiers}, p.~40, 2020.

\bibitem{weiland2018exploiting}
M.~Weiland, A.~Jackson, N.~Johnson, and M.~Parsons, ``Exploiting the
  performance benefits of storage class memory for hpc and hpda workflows,''
  \emph{Supercomputing Frontiers and Innovations}, vol.~5, no.~1, pp. 79--94,
  2018.

\bibitem{ren2019easycrash}
J.~Ren, K.~Wu, and D.~Li, ``Easycrash: Exploring non-volatility of non-volatile
  memory for high performance computing under failures,'' \emph{arXiv preprint
  arXiv:1906.10081}, 2019.

\bibitem{patil2019performance}
O.~Patil, L.~Ionkov, J.~Lee, F.~Mueller, and M.~Lang, ``Performance
  characterization of a dram-nvm hybrid memory architecture for hpc
  applications using intel optane dc persistent memory modules,'' in
  \emph{Proceedings of the International Symposium on Memory Systems}, 2019,
  pp. 288--303.

\bibitem{weiland2019early}
M.~Weiland, H.~Brunst, T.~Quintino, N.~Johnson, O.~Iffrig, S.~Smart, C.~Herold,
  A.~Bonanni, A.~Jackson, and M.~Parsons, ``An early evaluation of intel's
  optane dc persistent memory module and its impact on high-performance
  scientific applications,'' in \emph{Proceedings of the International
  Conference for High Performance Computing, Networking, Storage and Analysis},
  2019, pp. 1--19.

\bibitem{xiang2022characterizing}
L.~Xiang, X.~Zhao, J.~Rao, S.~Jiang, and H.~Jiang, ``Characterizing the
  performance of intel optane persistent memory: a close look at its on-dimm
  buffering,'' in \emph{Proceedings of the Seventeenth European Conference on
  Computer Systems}, 2022, pp. 488--505.

\bibitem{cost1}
V.~Kaplarevic, ``Intel optane memory vs ssds vs ram,''
  \url{https://phoenixnap.com/kb/optane-memory-vs-ssd-vs-ram }, December 24,
  2019, [Online].

\bibitem{cost2}
J.~Handy, ``Intel’s optane dimm price model,''
  \url{https://thememoryguy.com/intels-optane-dimm-price-model/ }, May 31,
  2019, [Online].

\bibitem{negevhpc}
R.~I. Park, ``Negevhpc project,'' \url{https://www.negevhpc.com}, 2019,
  [Online].

\end{thebibliography}

\documentclass[conference]{IEEEtran}
\usepackage[noadjust]{cite}
\usepackage{breakurl}
\usepackage{xurl}
\ifCLASSINFOpdf
\else
\fi

\DeclareRobustCommand*{\IEEEauthorrefmark}[1]{%
  \raisebox{0pt}[0pt][0pt]{\textsuperscript{\footnotesize\ensuremath{#1}}}}
  \usepackage[misc]{ifsym}
\newcommand{\corrAuthor}{$^{\textrm{\Letter}}$}
\begin{document}
\bstctlcite{IEEEexample:BSTcontrol}

\title{The Case for Non-Volatile RAM in Cloud HPCaaS \vspace{-0.5ex}}

\author{
\IEEEauthorblockN{Yehonatan Fridman\IEEEauthorrefmark{1,2}, Re'em Harel\IEEEauthorrefmark{1,3,4} and
Gal Oren\IEEEauthorrefmark{3,5\hspace{0.1cm}$\corrAuthor$}}\\
\IEEEauthorblockA{\IEEEauthorrefmark{1}Department of Computer Science, Ben-Gurion University of the Negev, Israel}
\IEEEauthorblockA{\IEEEauthorrefmark{2}Israel Atomic Energy Commission}
\IEEEauthorblockA{\IEEEauthorrefmark{3}Scientific Computing Center, Nuclear Research Center – Negev, Israel}
\IEEEauthorblockA{\IEEEauthorrefmark{4}Department of Physics, Nuclear Research Center – Negev, Israel}
\IEEEauthorblockA{\IEEEauthorrefmark{5}Department of Computer Science, Technion – Israel Institute of Technology, Israel}
{\tt\small fridyeh@post.bgu.ac.il, reemha@bgu.ac.il, galoren@cs.technion.ac.il}
\vspace{-0.5ex}
}

\IEEEtitleabstractindextext{%
\begin{abstract}
HPC as a service (HPCaaS) is a new way to expose HPC resources via cloud services. However, continued effort to port large-scale tightly coupled applications with high interprocessor
communication to multiple (and many) nodes synchronously,
as in on-premise supercomputers, is still far from satisfactory due to network latencies. As a consequence, in said cases, HPCaaS is recommended to be used with one or few instances. In this paper we take the claim that new piece of memory hardware, namely Non-Volatile RAM (NVRAM), can allow such computations to scale up to an order of magnitude with marginalized penalty in comparison to RAM. Moreover, we suggest that the introduction of NVRAM to HPCaaS can be cost-effective to the users and the suppliers in numerous forms.

\end{abstract}

\begin{IEEEkeywords}
NVRAM, Intel Optane\textsuperscript{TM}, HPC, Cloud, HPCaaS, Infiniband, DDR5
\end{IEEEkeywords}}
\maketitle

\IEEEdisplaynontitleabstractindextext

\IEEEpeerreviewmaketitle
\section{Supercomputing and the Cloud}

As the demand for more extensive calculations and simulations increases, huge amounts of data should be analyzed, manipulated, and stored. These needs ushered in a new Exascale era of stronger and more complex supercomputers \cite{heldens2020landscape, hoekstra2019multiscale, bernholdt2020survey}. This era was recently introduced with the first actual Exascale supercomputer – \textit{Frontier} \cite{schneider2022exascale}. \textit{Frontier} consists of more than 8 million compute cores that provide the peak computing power of 1.102 Exaflop/s.

Traditionally, said class of supercomputers can perform (almost) any kind of computation – from loosely to tightly coupled and from CFD simulations to deep-learning optimizations – with scalable performances compared to a single core or machine \cite{alexander2020exascale}. This scalability is mandatory for R\&D breakthroughs, as exact and more detailed calculations are the key to better science, and they tend to be latency-sensitive \cite{murphy2007effects}. Thus, the network architecture choice in modern supercomputers is of high speed, wide bandwidth, low latency, RDMA transmission enabled network, namely InfiniBand \cite{pfister2001introduction, grun2010introduction}, that is used to construct a complete system without (almost) any interference \cite{ajima2022optical, lu2022survey}. With switch latencies close to 100$ns$, and end-to-end latency of less than 1$\mu$s (end-point to end-point), InfiniBand interconnect achieves comparable performances to shared memory communication for high-performance and data-intensive applications. In scaling measurements of \textit{Summit} supercomputer, not long ago the world's fastest supercomputer, latency experiments from/to all 4,608 nodes showed a latency of 1.3$\mu$s for 0B messages and 3.54$\mu$s for 4KB messages with Mellanox EDR InfiniBand \cite{vergara2019scaling}.

Alongside the evolution of ‘classic’ supercomputing, a new architecture for HPC was raised – The Cloud, specifically HPC as a service (HPCaaS), which exposes HPC resources via cloud services \cite{parashar2013cloud}. The recent significant developments in the cloud allows better compute nodes than those in supercomputers, with denser computing power and much-improved access and elasticity \cite{aws2, aws3, gcp4}. As such, many small to low-medium scale calculations that usually needed access to supercomputers to be executed now can easily port to cloud services \cite{walkup2022best}. However, continued efforts to port large-scale tightly coupled applications with high interprocessor communication to multiple (and many) nodes synchronously, as in on-premise supercomputer, is still far from satisfactory and currently measured with a few hundred cores before scalability crashes \cite{fernandez2021evaluation}. This scalability issue results from the fabric (Ethernet, in this way or the other) with latency as high as one to two orders of magnitude than Infiniband on-premise supercomputers, even with dedicated optimizations \cite{fernandez2021evaluation, aws1, gcp1, gcp2, gcp3, phanekham2020measuring}. As a result, the cloud is usually recommended to be used in such scenarios with just one powerful instance (or a few), with as much computing power, memory and storage \cite{walkup2022best,aws6, zivanovic2016large}. For example, the new AWS Hpc6a.48xlarge contains 96 cores with 384GB of RAM \cite{aws2}, and the majority of cloud services already make available local SSD storage through SCSI and even NVMe interfaces \cite{kashyap2022nvme,guz2018performance,gupta2011evaluation, gcp5, aws4}. 

From the business perspective, the economic model of the public cloud, that much prefers elasticity over connectivity speed, can not support Infiniband networking efficiently and cost-effectively, let alone supplying a large amount of resources simultaneously, spontaneously and instantly \cite{netto2018hpc}. A unique exception that does not prove the rule is a supercomputer that was fully ported to the cloud: In Nov 2021, and for the first and only time since then, new supercomputers (mainly No. 10 on the TOP500 list \cite{top500}, the \textit{Voyager-EUS2}) were introduced entirely to one of the major cloud suppliers. As such, \textit{Voyager-EUS2} enjoys the two technologies’ benefits and disregards their weaknesses, except the most critical one in the cloud, the elasticity, that is bounded to the system resources.

In a seminal paper from 2009, Napper and Bientinesi wondered "Can Cloud Computing Reach The TOP500?" \cite{napper2009can}, and their concerns remain valid. Specifically, their observation that the flop/s obtained per dollar spent decrease exponentially with increasing computing cores and correspondingly, the cost for solving a linear system increases exponentially with the problem size -- is still very much in contrast to existing scalable supercomputers. "If cloud computing vendors are serious about targeting the HPC market," they stated, "different cost models must be explored. An obvious first step would be to offer better interconnects or nodes provisioned with more physical memory to overcome the slower network" \cite{napper2009can}. 

\section{Non-Volatile RAM in HPC Use cases}
Three primary resources of HPC (node-wise, excluding communication) are computing power, main memory and storage. Each component develops differently, which creates significant gaps that impair the scalability and performance of these systems. Large capacity of main memory is required to execute large problems, while its low latency is required to constantly feed the computing units with data. While computing power has increased exponentially over the years (mainly due to integrating heterogeneous hardware like CPUs and GPUs) \cite{denning2016exponential,top500}, DRAM scales slowly and presents no significant improvement in latency or bandwidth: The world's first DDR5 DRAM chip was officially launched in 2020 \cite{jing2020cache}, a decade after the DDR4 was introduced in 2011; and it has about the same latency as DDR4, just four times the density, and is only targeted to double the bandwidth. For example, in the flagship\textit{ Aurora} supercomputer \cite{aurora}, each compute node will hold the top density possible of 16 dies times 64GB per die, meaning $\sim$1TB. The DRAM is also expensive relative to capacity, creating a significant gap with the increasing demand for computational power.
Worse, the storage performances of HDDs and SSDs are extremely low in comparison to the DRAM, and exhibit no significant improvements in technology over the years. There are still 2-3 orders of magnitude between the performances of DRAM and storage devices \cite{luttgau2018survey}, and as such they create a huge bottleneck to scientific applications \cite{harrod2012journey}. However, DRAM is significantly more expensive than standard storage devices~\cite{luttgau2018survey} and its power consumption is higher, due to the cost of frequent refresh operations~\cite{fan2001memory}. Thus, building cost-effective, resilient, fast and power-efficient systems with large memory and storage spaces is a major challenge faced by the HPC community~\cite{oren2016memory,bergman2008exascale}.

A new building block in supercomputers is \textit{non-volatile RAM} (\textit{NVRAM}). Next-generation supercomputers are expected to integrate NVRAM devices. For example, \textit{Aurora}~\cite{Argonne1} is planned to integrate the emerging Intel Optane™ NVRAM. NVRAM provides better density and energy efficiency than DRAM
while providing DRAM-like performance: NVRAM comes in 128GB, 256GB and 512GB capacities (vastly larger than DRAM cards) \cite{intel1} and its read/write latencies are 400\% higher than those of DRAM, while the read/write bandwidths were 37\% and 8\% of those of DRAM, respectively \cite{hirofuchi2020prompt,hirofuchi2019preliminary}.
When configured in App-Direct mode these devices are byte-addressable and can be used by processes as NVRAM, while when configured in Memory or Flat mode these devices provide an extension to the volatile memory pool of the application, enabling the execution of much larger problems in a small number of compute nodes \cite{patil2021nvm}. 

Prior work on the use of NVRAM in HPC environments was focused on three main direct use-cases: 
(1) as fast storage for diagnostics~\cite{fridman2021assessing,hennecke2020daos,weiland2018exploiting}, 
(2) as fast persistence area for checkpointing~\cite{fridman2021assessing,ren2019easycrash}, and 
(3) as memory expansion to enable larger memory scientific workloads~\cite{fridman2021assessing,patil2019performance,weiland2019early,patil2021nvm}. 
The two former use-cases rely on the non-volatility of NVRAM for fast storage as a substitute for standard storage mediums~\cite{weiland2018exploiting,hennecke2020daos,fridman2021assessing}. 
The third use-case investigates the benefits of NVRAM in Memory mode, using its large capacity to enhance HPC capabilities~\cite{patil2019performance,patil2021nvm,fridman2021assessing}. 
In particular, a comprehensive investigation~\cite{fridman2021assessing} studies the main use-cases of Intel Optane™ NVRAM in HPC to improve scientific applications’ performances and fault tolerance.

\section{The Case for Non-Volatile RAM in HPCaaS}

It was shown that various HPC applications can benefit from scaling-in applications into fewer nodes by integrating NVRAM as memory expansion --- in terms of energy efficiency and performance \cite{patil2021nvm}. Patil et al. \cite{patil2021nvm} investigated the trade-off between performance degradation due to the higher latency of NVRAM and the savings in intra-node communication overheads when executing various scientific applications on a single NVRAM node (which contains DRAM and NVRAM) rather than on 4 DRAM-only nodes. This investigation assumed an Infiniband connection between the 4 DRAM-only nodes, and found that the shared memory buffer communication on the NVRAM node has up to half the latency and 3x the bandwidth than the Infiniband-connected DRAM-only nodes. Patil inferred that compute-bound applications, which utilize cache locality well, are able to run larger problem sizes on fewer compute nodes with a DRAM-NVRAM hybrid memory system, experiencing minimal performance degradation while providing significant cost and energy savings.

As a consequence of previous conclusions by Patil et al. \cite{patil2021nvm}, and given the Ethernet-based cloud connectivity, we hereby take the claim that NVRAM is an essential, and even crucial component in future cloud HPCaas instances, when considering memory and I/O bounded applications in general, or small to mid-range tightly-coupled computations in particular. We notice that even in current NVRAM capacity, it is possible -- with only a few factors of performance degradation in comparison to RAM \cite{fridman2021assessing, weiland2019early, xiang2022characterizing} -- to achieve up to an order of magnitude more byte-addressable memory, with much lower costs \cite{cost1, cost2}. In weak-scaled computations, that rely on RAM extensions, the presence of such memory can save not only the need to extend to other instances/nodes (and by that to reduce dramatic latency overheads), but also to significantly reduce the costs -- either for the clients (by requesting less resources and achieving more scalability) or the cloud supplier (with energy and scheduling savings). Furthermore, in applications that heavily rely on I/O, either for diagnostics or even for recoverability, the local NVRAM can serve as a reliable, vast and fast storage and memory combined, reducing the need to extend external distributed file systems over high latency communication.

\clearpage
\section*{Acknowledgment}
This work was supported by Pazy grant 226/20, ISF grant 1425/22, and the Lynn and William Frankel Center for Computer Science. Computational support was provided by the NegevHPC project~\cite{negevhpc}. 

\bibliographystyle{IEEEtran}
\bibliography{sample-base.bib}

\end{document}